OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = 18, b = 26, c = ?

OpenStudy (akashdeepdeb):

Pythagorus' theorem = \(a^2 + b^2 = c^2 \implies c = \sqrt{a^2 + b^2}\)

OpenStudy (anonymous):

still confused lol

OpenStudy (anonymous):

just think what answer u think it is always listen to ur gut

OpenStudy (phi):

can you write down pythagoras formula?